The histogram displays the number of 2007 births among U.S. women ages 1010 to 5050. Each bin represents an interval of two years, and the height of each bin represents the frequency with which the data fall within that interval.

Births by mothers age (2007) 483486 455 500 463 427 403 3 400 375 330 300 265 205 200 176 98 100 45 17 4 0 10 20 30 40 50 Mothers age (in years)

Use the provided histogram to answer questions about the number of births for each age group.


1) How many births occurred among women over the age of 4040?

number of births:

___________________________

2) What percentage of births occurred among women between the ages of 3838 and 4040? Give your answer to one decimal place.

percentage of births:

___________________________

3) What does the histogram tell you about the ages at which women gave birth in 2007?

A)The distribution of ages at which women gave birth is roughly symmetric.

B)Four women between the ages of 4444 and 4646 gave birth in 2007.

C)No woman under the age of 1616 gave birth in 2007.

D)Women between the ages of 2626 and 2828 gave birth more often than any other age group.
Answer #1

1) over age of 40 birth occurred = 45+17+4 =      66

so, answer is 66000
                  
2) total birth occurred between age 38 and 40 = 98                  
                  
total births= 66+205+375+427+455+483+486+463+403+330+265+176+98+45+17+4 =4298              
                  
so, percentage=98/4298*100=   2.280%

3)

D)Women between the ages of 26 and 28 gave birth more often than any other age group.